OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = n(6n^2-3n-1)/2

OpenStudy (anonymous):

using http://openstudy.com/study#/updates/504b9d7ce4b0985a7a58b494 I understand everything until

OpenStudy (anonymous):

I don't get how the last form is "precisely what we were supposed to get" and how it proves it true

OpenStudy (usukidoll):

hmmm... first we need the basis then when we have k and then k+1 have to manipulate the left to equal the right ..

OpenStudy (michele_laino):

for n= 1, we have: left side = 1^2 = 1 right side= 1*(6-3-1)/2= 1 so our proposition is true for n=1

OpenStudy (usukidoll):

ok so we have the basis proven true.. next we need to have n = k

OpenStudy (michele_laino):

in general, the mathematical induction principle requests another checking, for example n=2

OpenStudy (anonymous):

Ok I understand why and how to do that

OpenStudy (michele_laino):

so, for n=2, we have: left side = 1+ 4^2=1+16=17 right side = 2*(24-6-1)/2= 17

OpenStudy (anonymous):

but what about n = k + 1? how is the simplified form of n = k+1 proving it true???

OpenStudy (usukidoll):

first we need n = k and then we go to n = k+1 . That guy went forward to n+1

OpenStudy (michele_laino):

when we have a proposition which depends on a natural number, say n, namely, we have: P(n), then if P(1) is true and P(n+1) is true when P(n) is true, then we can state, that P(n) is true for all natural numbers, namely: \[\Large \forall n \in \mathbb{N}\]

OpenStudy (usukidoll):

\[1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = \frac{n(6n^2-3n-1)}{2}\] for n = k \[\[1^2 + 4^2 + 7^2 + ... + (3k - 2)^2 = \frac{k(6k^2-3k-1)}{2}\]\] for n = k+1 \[\[1^2 + 4^2 + 7^2 + ... + (3(k+1) - 2)^2 = \frac{(k+1)(6(k+1)^2-3(k+1)-1)}{2}\]\]

OpenStudy (usukidoll):

so to simplify this mess a bit ... sorry this is long! \[\[1^2 + 4^2 + 7^2 + ... + (3k+1)^2 = \frac{(k+1)(6(k+1)^2-3(k+1)-1)}{2} \] \[\[1^2 + 4^2 + 7^2 + ... + (3k+1)^2 = \frac{(k+1)(6(k+1)^2-3k-4)}{2} \] \[\[1^2 + 4^2 + 7^2 + ... + (3k+1)^2 = \frac{(k+1)(6(k^2+2k+1)-3k-4)}{2} \] \[\[1^2 + 4^2 + 7^2 + ... + (3k+1)^2 = \frac{(k+1)(6k^2+12k+6)-3k-4)}{2} \] \[\[1^2 + 4^2 + 7^2 + ... + (3k+1)^2 = \frac{(k+1)(6k^2+12k-3k-4+6)}{2} \] \[\[1^2 + 4^2 + 7^2 + ... + (3k+1)^2 = \frac{(k+1)(6k^2+9k+2)}{2} \]

OpenStudy (usukidoll):

hmmm then manipulate the left... hold on this is in my book I need to refresh my memory on it.. I remember substituting that 1^2+4^2+7^2

OpenStudy (usukidoll):

I know that our goal is to have the left = right ... I just have to see one of the examples in my book again.

OpenStudy (usukidoll):

a proposition was used... For all n in N  then the entire 1^2 4^2 7^2 is substituted . we do know that \[(3k+1)^2 = (3k+1)(3k+1) = 9k^2+6k+2 \]